OpenStudy (anonymous):

find the inverse function f(x)= 4-x over 2x+3

OpenStudy (zzr0ck3r):

y = (4-x)/(2x+3) we want to solve for x y(2x+3) = 4-x 2xy+3y = 4-x 2xy+x = 4-3y x(2y+1) = 4-3y x = (4-3y)/(2y+1) now switch the x and y y = (4-3x)/(2x+1)

OpenStudy (zzr0ck3r):

so the inverse is f^(-1)(x) = (4-3x)/(2x+1)

OpenStudy (zzr0ck3r):

make sense?

OpenStudy (zzr0ck3r):

@glares25 ? i see you bumped the question but i have answered it for you. is there anything you dont understand?

OpenStudy (anonymous):

no i think your right , i just confused with my own answer thank you so much

OpenStudy (zzr0ck3r):

let me know if anythng does not make sense. or yif you want to show me what you did we can see whats going on there.

OpenStudy (anonymous):

thank you, my answer is -4 - x / 2x-3 but i think its nonsense

OpenStudy (zzr0ck3r):

can you show me the steps you took to get there?

OpenStudy (zzr0ck3r):

it seems to close to be nonsense prob a small error.

OpenStudy (anonymous):

will you explain me x(2y+1) = 4-3y, i dont know how u get the 1 there. thank you

OpenStudy (zzr0ck3r):

ok the line before that was 2xy+x = 4-3y so I factored out an x on the left hand side and get x(2y+1) = 4-3y

OpenStudy (zzr0ck3r):

a(b+c) = ab+ac

OpenStudy (anonymous):

ahhh ok thank you

OpenStudy (zzr0ck3r):

np